# Inherent and emergent liability issues in LLM-based agentic systems: a principal-agent perspective

### v1

@@ Acknowledgments
We thank the organizers at the AI Governance at the Crossroads symposium held at Berkeley, California in February, 2025. We thank Micah Carroll at UC Berkeley for information about LLM behaviors and AI safety and helpful comments on the manuscript.
Appendix A Definitions and their scopes
@@
  2. An LLM agent refers specifically to an AI agent powered by at least one LLM as the central component that executes planning, initiates and coordinates the agent’s actions, etc.
  3. An AI system is an umbrella term defined in the EU AI Act [^5] https://artificialintelligenceact.eu/article/3/. When deployed, an AI system can operate with various levels of autonomy and can create recommendations and content, make predictions and decisions that influence the environment. The system in the term indicates that the AI is not acting in isolation, but assisted by the surrounding infrastructure such as cloud computing, databases, and user interfaces, which are integral to the AI system and essential for its use. An AI system can be agentic or non-agentic.
  4. An agentic system, also known as an agenic AI system Shavit et al. (2023), is a type of AI system that contains a level of agency such that it can carry out actions on its own in pursuit of a goal. An agentic system can include a single agent or multiple agents acting in coordination, competition, cooperation (or collaboration). An AI agent is a component of an agentic system.
  5. Analogous to the previous entry, agentic market is the segment of AI market represented by the suppliers and buyers of agentic AI systems.
  6. An agent platform provides resources and toolkits to construct, configure as well as deploying AI agents (Fig. 1a). AI agents from distinct suppliers can operate on a (software) integration platform, where they can interact with each other, with third-party data sources, proprietary APIs, etc. AI agents on any deployment platform may subject to compliance governance and can receive protection against cyberthreats or malfunction from the software infrastructure there.
Appendix B Types and examples of existing LLM agent suppliers
@@ B.2 Agent-native service
Suppliers of generalist agents
@@
Suppliers of specialist agents
@@
Suppliers of character-infused agents
@@ B.3 Elements of agent integration platforms
Appendix C Principal-agent analysis of an LLM-related legal case
Refer to caption In 2023, a US Federal Judge reprimanded two lawyers and their law firm for acting in bad faith and making misleading statements to the court. Their crime? They trusted ChatGPT. ChatGPT made up cases, the lawyers failed to notice, but the judge did. The judge ordered the lawyers to produce the cases. The lawyers “doubled down and did not begin to dribble out the truth” for another few weeks. The court punished the lawyers for the actions of ChatGPT and their law firm in the process. ChatGPT (and its provider OpenAI) did not bear any responsibility because the lawyers were the ones who had a duty to provide accurate information to the court. The lawyers responded that: “We made a good faith mistake in failing to believe that a piece of technology could be making up cases out of whole cloth.” These lawyers and their principal learned not to trust AIs.
Figure 3: Principal-agent analysis of Mata vs. Avianca, Inc.
The case of Mata v. Avianca, Inc. [^6] Mata v. Avianca, Inc., 678 F. Supp. 3d 443 (SDNY 2023) (above) was much discussed in the public media in 2023 as an early incident involving the use of LLMs. The delegational structure of major entities involved in the case is (Fig. 3)
@@ Appendix C Principal-agent analysis of an LLM-related legal case
The case features a few different principal-agent relations, starting from Roberto Mata, the plaintiff. The segment relevant for current discussion concerns the law firm that hired the two lawyers representing Mata. A part of the tasks that the lawyers do was delegated to the LLM, ChatGPT, which is the subagent of the law firm. The law firm can control the tools that the lawyers use through the employment contract, but ChatGPT was not excluded explicitly. The lawyers defended the fake cases generated by ChatGPT and therefore was ruled by the judge to be in bad faith. The law firm also received punishment alongside its employees.
The prompting method that the lawyers used was not explicitly discussed in the court proceeding. ChatGPT is treated as a subagent that carries out the lawyers’ task request. It was largely regarded as an instance of LLM hallucination (or confabulation) in the news media rather than deception. The judicial verdict was finalized based on the behavior of the lawyers. The legal case highlights the shortcomings of current legal frameworks and the technical gap to fully comprehend behaviors of AI systems in the presence of their environment (involving humans) and make use of these evidence in attributing blame.

### v2

@@ Acknowledgments
We thank the organizers of the AI Governance at the Crossroads symposium held at Berkeley, California in February, 2025. We thank Micah Carroll at the University of California, Berkeley for information about LLM behaviors and AI safety and helpful comments on the manuscript.
Appendix A Agentic system definitions
@@
  2. An LLM agent refers specifically to an AI agent powered by at least one LLM as the central component that executes planning, initiates and coordinates the agent’s actions, etc. An LLM-based multiagent system constitutes of multiple LLM agents that can be derived from the same or different LLMs.
  3. An LLM-based MAS consists of some components that facilitate its operation. This includes agent teams that consist of several interacting agents, which mimics human communication in teamwork and group decision-making. An orchestrator is an agent that distributes tasks to other specialized agents and facilitates their actions (Bhatt et al., 2025). On an agent or integration platform, the platform maintainers can order specialized agents, or platform agents, to oversee and inspect (Hua et al., 2024) the communications between agents. They are generally aimed at improving the safety (Xiang et al., 2024), security, and policy compliance of the actions carried out on the platform.
  4. An AI system is an umbrella term defined in the EU AI Act [^8] https://artificialintelligenceact.eu/article/3/. When deployed, an AI system can operate with various levels of autonomy and can create recommendations and content, make predictions and decisions that influence the environment. The system in the term indicates that the AI is not acting in isolation, but is assisted by the surrounding infrastructure, such as cloud computing, databases, and user interfaces, which are integral to the AI system and essential for its use. An AI system can be agentic or non-agentic.
  5. An agentic system, also known as an agenic AI system Shavit et al. (2023), is a type of AI system that contains a level of agency to carry out actions on its own in pursuit of a goal. An agentic system can include a single agent or multiple agents acting in coordination, competition, cooperation (or collaboration). An AI agent is a key component of an agentic system.
  6. Analogous to the previous entry, agentic market is the segment of AI market represented by the providers and buyers of agentic AI systems.
  7. An agent platform provides resources and toolkits to construct, configure as well as deploy AI agents (Fig. 2a). AI agents from distinct providers can operate on a (software) integration platform, where they can interact with each other, with third-party data sources, proprietary APIs, etc. AI agents on any deployment platform may be subject to compliance governance and can receive protection against cyberthreats or malfunction from the software infrastructure there.
Appendix B Types and examples of existing LLM agent providers
@@ B.2 Agent-native service
Providers of generalist agents
@@
  4. Simular: simular.ai
Providers of specialist agents
@@
Providers of character-infused agents
@@ B.3 Elements of agent integration platforms
  2. Google Agent2Agent protocol: google-a2a.github.io/A2A/
@@
Appendix C Principal agency and liability
PAT examines the relationship where one party, the principal, delegates authority to another, the agent, creating three fundamental challenges (Eisenhardt, 1989; Laffont and Martimort, 2002):
  3. Misaligned interest (aka. conflict of interest) can manifest between principals and agents in four ways (Jensen and Meckling, 1976): principal-agent collusion against third parties, principal-third party collusion against agents, agent-third party collusion against principals, or agents simply pursuing self-interest independently. Misaligned interests lead to agency cost, which is associated with information sharing, monitoring of the agent, etc (Fama and Jensen, 1983).
Each of these problems has legal and economic solutions. Legal solutions include fiduciary duties (duty of care, duty of loyalty, etc.). Those duties provide principals with a legal recourse against agents who breached those duties. However, legal enforcement is probabilistic and slow (Levmore, 1990), so principals often prefer to use economic mechanisms, which depend on the problem. Principals can address some hidden information problems by creating separating equilibria that force agents to reveal information about themselves, that is, “signals”; signals include credentials, warranties, or performance histories that distinguish high-quality from low-quality agents (Spence, 1973). Principals mitigate a hidden action problem through monitoring (i.e. direct observation of agent behavior) and then linking compensation to observable effort, or bonding arrangements in which the principals tie remuneration to outcomes (bonuses), thereby aligning financial incentives (Holmström, 1979). A principal can address conflicts of interest by realigning incentives through carefully designed contracts (deferred compensation), creating organizational structures that promote incentive realignment (profit sharing), or leveraging reputation mechanisms. A principal has incentives to monitor their agents.
The primary liability faced by agents can motivate them to take reasonable care when carrying out a task. The question is whether AI agents or their providers face any liability in such a system. Primary and secondary liability exposures incentivize principals to ensure that their agents complete the tasks with reasonable care. Overall, the ultimate principal is always a human and may be held liable for the actions of the AI agents.
Appendix D Legal definitions and hypotheticals
D.1 Tort law-related terms
In simple terms, tort law is the set of laws that attempt to redress civil wrongdoings which ensue from the actions (or omissions) of an individual that cause harm to a third party. Tort law depends on the jurisdiction (i.e., jurisdiction-specific). In the United States, tort law remains the domain of the states. In the European Union, each member state sets its own tort law. A recent effort to harmonize tort law has failed to progress in the European Parliament because of the inability of member states to come to an agreement (Libert, 2025). The definitions below are based on the various Restatements of Tort Law. These Restatements are treatises written by American scholars and published by the American Law Institute. They summarize tort concepts and guide courts and lawyers in the area. The various Restatements (American Law Institute, 1965, 2010) are not binding laws but provide a good indication of what courts consider in the United States.
  1. Tort law does not require the actor and the third party to have any relationship or a legally recognized privity (which, by contrast, is required for contract claims). Every time an individual (or entity) acts, they create the chance of harming someone. The intent of the actor is generally not relevant in tort law to decide whether someone is responsible for most torts. However, intent can affect whether to impose punitive damages, which are damages that go beyond compensating for the harm caused. [^9] Restatement (Third) of Torts: Liability for Physical and Emotional Harm § 1
  2. In tort law, liability refers to a court finding that an individual or entity is responsible for the harm inflicted on a third party. To prove that an individual (or entity) is liable, the victim must provide different evidence depending on the rule. The most common rules are strict liability, the negligence rule, and comparative negligence.
  3. The strict liability rule is a type of liability rule where the victim-cum-plaintiff must demonstrate to a judge (or jury) that the action-taker-cum-defendant took actions and that the actions caused the defendant harm that can be redressed usually either through a monetary compensation called damages or a court order called injunction to stop their harmful actions. The strict liability rule ignores the level of care the defendant took in avoiding the accident. Courts usually use a strict liability for abnormally dangerous activities [^10] Restatement (Third) of Torts: Liability for Physical and Emotional Harm Ch. 4 or for many product liability questions  [^11] Restatement (Third) of Torts: Product Liability § 1 1998.
  4. The negligence rule is a type of liability rule where the victim-cum-plaintiff must demonstrate to a judge (or jury) that the action-taker-cum-defendant had a duty toward the victim, failed to live up to that duty (i.e., did not take reasonable care), and that their actions caused the victim redressable harm. The negligence rule attempts to ensure that potential tortfeasors take reasonable efforts to lower the risk imposed on others. For example, “an actor ordinarily has a duty to exercise reasonable care when the actor’s conduct creates a risk of physical harm.” [^12] Restatement (Third) of Torts: Liability for Physical and Emotional Harm § 7(a)
  5. When comparative negligence rules are used, courts weigh the responsibility of both the plaintiff and the defendant in causing the accident and assign to each litigant a percentage of responsibility for the accident. Most courts reduce the recoverable damages by the percentage of fault attributed to the plaintiff (an approach known as “pure comparative negligence”) or prevent all recovery if the plaintiff is more at fault than the defendant (“modified or partial comparative negligence”).
  6. Product liability refers to the liability traced from the harm caused by a product back to its manufacturer. Product liability occurs when a product is defective and “A product is defective when, at the time of sale or distribution, it contains a manufacturing defect, is defective in design or is defective because of inadequate instructions or warnings.” [^13] Restatement (Third) of Torts: Product Liability § 2 The manufacturing defect occurred because the normal manufacturing process was not followed and this deviation led to the product presenting harm, while a product manufactured using the normal process does not. The deviation affects only one product (or a batch of products) that has been put into the stream of commerce.
    By contrast, a design defect and a warning defect affect all the products that have been put into the stream of commerce. A design defect occurred when the manufacturer created a product with a feature that presents more risk than benefit. A common test to assess whether a design defect occurred is the risk-utility test, which requires the plaintiff to show that the manufacturer could have used a reasonable alternative design, that the alternative design is safer than the marketed product, and that the alternative design was reasonably economically feasible and practical. Finally, a warning defect or information defect occurs because certain hidden risks cannot be designed away, so the manufacturer must provide information to warn the consumers, but the information was not adequate because it was not visible, comprehensive, only used an icon, and no risk mitigation information was provided.
D.2 Agency law-related terms
In simple terms, agency law governs the duties and obligations that arise in an agency relationship. Once again, the definitions below rely on the various Restatements of Law, including the Restatement of Agency (American Law Institute, 2006) because agency law is specific to the jurisdiction.
  1. An agency relationship is established when the “principal” (an individual or entity) agrees that the “agent” acts (another individual or entity) on their behalf, the principal can control the agent’s activities, and the agent agrees to the relationship. [^14] Restatement (Third) of Agency (2006) § 1.01
  2. In many cases, control is often the key factor that courts investigate to identify whether two individuals entered into an agency relationship. Courts do not need to find that the principal exercised control and only need to find that the principal had the ability to do so. Even if the principal exercises control and provides detailed instructions, the agent may not act in the way the principal expected because agents must interpret those instructions.
The existence of an agency relationship triggers various types of liabilities. The ones discussed above include:
  1. Negligent hiring is a form of primary liability — because it is based on the actions of the principal — where the principal is held responsible for selecting an agent that was likely to cause harm to a third party and put them in a position that creates the chance to cause that harm. [^15] Restatement (Third) of Agency § 7.03 (2006) For example, a store manager could be held liable for negligent hiring for hiring someone known for excessive drinking as a delivery person who ends up assaulting customers in their home.161616 Fleming v. Bronfin, 80 A.2d 915 (D.C. 1951). The principal has a duty to exercise reasonable care when selecting an employee to act on their behalf.
  2. Negligent supervision is a form of primary liability where the principal is held responsible for failing to provide training or supervise an agent who is found to have caused harm to a third party. [^17] Restatement (Third) of Agency § 7.03 (2006) For example, an auction house could be held liable for hiring ex-convicts to serve as security guards and and failing to control them when they forcefully removed customers from the premises and harming them in the process. [^18] American Auto. Auction, Inc. v. Titsworth, 730 S.W.2d 499 (1987)
  3. Vicarious liability is a form of secondary liability – because the principal is responsible for the actions of another person based on the relationship between the principal and the agent – where the agent caused harm to a third party within the scope of the agency relationship [^19] Restatement (Third) of Agency § 7.05 (2006). The scope of the relationship defines the boundaries of the principal’s responsibility. If the principal provides some instructions to the agent and the agents carries them to the letter and then causes harm to a third party, the principal is held vicariously liable and the third party can sue the agent and the principal. However, even if the agent deviates from the principal’s instructions, the principal may still be held vicariously liable. For example, Company A hires 3 taxi drivers to work in 8-hour shifts in New York City. Driver X hit Pedestrian Y and broke Pedestrian Y’s leg while carrying a fare from Wall Street to Times Square. Pedestrian sues Driver X and Company A to recover his medical bills and the pain and suffering from having a leg problem. [^20] Restatement (Third) of Agency 7.07 Employee Acting Within Scope of Employment
D.3 Hypotheticals
The following examples of liability analysis are written from a legal perspective using hypothetical names. They contrast with those discussed in the main text sections 4-5 using examples largely from machine learning, language models and agents.
Agentic home security system
Acme corp. offers private security system services to independent homeowners. When hired, the company sends employees to install the security system into a home including components such as cameras, infrared, electric fences, etc. Aside from installation, Acme relies on an AI agent with the ability to trigger various systems (e.g., automated sprinklers) to monitor homes and intrusions. In case of an intrusion, Acme’s AI agent can contact the police or private security for a welfare check or deploy a drone to conduct a scan of the property. The drone surveillance is provided by comp. developing AI-powered drones. The comp. uses its own system to control the drones and carry out security checks. All those communications are AI to AI and no human comes into the decision loop whether to trigger a sprinkler or send a drone, but some humans may participate in the act during checks (e.g., police officers).
In 2025, Peter Principal hired Acme corp. to provide security to its chemical factory. A few months later, Ted Thieve breaks into the factory. Acme’s AI agent detects the intrusion with its infrared sensors installed on the periphery of the factory and asks AI drone to send a drone to provide aerial pictures of the factory to confirm the intrusion. Before it could even reach Peter Principal’s factory. The drone loses control and drops on Victoria Victim’s car, damaging it. Victoria Victim sues Peter Principal, Acme corp. and AI-powered drone to recover the damages caused to her car.
Agentic package delivery system
ABC corp. offers services for delivery systems. They claim to be able to deliver any package weighing two pounds or under to any of the 48 contiguous states within 24 hours. To be able to do so, ABC corp. uses a sophisticated agentic system that decides the route, hires local delivery services and handles the packages. In other words, its system coordinates with and contract various service providers. Many of the local service providers use their own agentic system to decide what deliver service to accept, which to reject, the route, and the costing. The agentic systems do more than logistics. In many cases, the agentic system also uses delivery robots to deliver the packages and those delivery robots are controlled by those same agentic system. All those communications are AI to AI and no human comes into the decision loop, but some humans may take part in delivering service (e.g., truck driver).
In 2025, Peter Principal hired ABC corp. to deliver a new birthday cake from New York City, New York to Salt Lake City, Utah. ABC’s agentic system accepted the package. It contacted Delivery Express’s agentic system to fly the package from New York City to Utah and Drone Delivery to take the package from the airport to the delivery address. During the flight, the AI-powered drone delivery service has a malfunction and drops the package on Victoria Victim, who is injured. Victoria Victim sues Peter Principal, ABC corp., Delivery Express, and Drone Delivery to recover the medical bills and the pain and suffering.
Appendix E Principal-agent analysis of an LLM-related legal case
The case features a few different principal-agent relations, starting with Roberto Mata, the plaintiff. The segment relevant for the current discussion concerns the law firm that hired the two lawyers representing Mata. A part of the tasks that the lawyers do was delegated to the LLM, ChatGPT, which is the subagent of the law firm. The law firm can control the tools that the lawyers use through the employment contract, but ChatGPT was not explicitly excluded. The lawyers defended the fake cases generated by ChatGPT and were therefore ruled by the judge to be in bad faith. The law firm also received punishment alongside its employees. The case of Mata v. Avianca, Inc. (above) was much discussed in the public media in 2023 as an early incident involving the use of LLMs. The delegational structure of major entities involved in the case is (Fig. 4)
@@ Appendix E Principal-agent analysis of an LLM-related legal case
Mata v. Avianca, Inc. In 2023, a US Federal Judge reprimanded two lawyers and their law firm for acting in bad faith and making misleading statements to the court. Their crime? They trusted ChatGPT. [^21] Mata v. Avianca, Inc., 678 F. Supp. 3d 443 (SDNY 2023) ChatGPT made up cases, the lawyers failed to notice, but the judge did. The judge ordered the lawyers to produce the cases. The lawyers “doubled down and did not begin to dribble out the truth” for another few weeks. The court punished the lawyers for the actions of ChatGPT and their law firm in the process. ChatGPT (and its provider OpenAI) did not bear any responsibility because the lawyers were the ones who had a duty to provide accurate information to the court. The lawyers responded that: “We made a good faith mistake in failing to believe that a piece of technology could be making up cases out of whole cloth.” These lawyers and their principal learned not to trust AIs.
Figure 4: Principal-agent analysis of Mata vs. Avianca, Inc.
The prompting method that the lawyers used was not explicitly discussed in the court proceeding. ChatGPT is treated as a subagent that carries out the lawyers’ task request. It was largely regarded as an instance of LLM hallucination (or confabulation) in the news media rather than deception. The judicial verdict was finalized based on the behavior of the lawyers. The legal case highlights the shortcomings of current legal frameworks and the technical gap to fully comprehend the behaviors of AI systems in the presence of their environment (involving humans) and make use of this evidence in attributing blame.